solve system by addition, then state Y value
x+y=28
x-y=12

Answer by checkley71(8403)

X+Y=28
X-Y=12
---------------------------------
2X=40
X=40/2
X=20 ANSWER.
20+Y=28
Y=28-20
Y=8 ANSWER.
PROOF
20-8=12
12=12